OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. Please show all of your work. Then graph the equation and submit your graph through the dropbox. 9y – 3x = -6

OpenStudy (radar):

(2,0) and (0,-2/3)

OpenStudy (anonymous):

heree is a hint, you use the substitution property so you choose which one you want to get alone first

OpenStudy (anonymous):

This is in standard form. To find the x intercept 9(0) - 3x = -6 -3x = -6 -3x/-3 = -6/-3 x = 2 x intercept = (2,0) Find the y intercept 9y - 3(0) = -6 9y = -6 9y/9 = -6/9 y = -6/9 or -2/3 y intercept = (0, -2/3)